OpenStudy (anonymous):

(Will give medal) Which is an equation in point-slope form for the given point and slope? Point: (1, 9); Slope: 5 A. y – 1 = 5(x + 9) B. y – 9 = 5(x + 1) C. y – 9 = 5(x – 1) D. y + 9 = 5(x – 1)

OpenStudy (anonymous):

plug the info into y-y1=m(x-x1) y1 is the y coordinate x1 is the x coordinate m is the slope y-9=5(x-1)

OpenStudy (anonymous):

Thanks so much!